6 consecutive whole numbers on the number line. The number in
position two is a multiple of 11 and the number in position six is a multiple of 10. Find the six number.

Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

number in position two = n
number in position six = n + 4

n = 11

6 consecutive numbers in order: 10, 11, 12, 13, 14, 15

Answer by ikleyn(52856)

.

            The solution by  @CubeyThePenguin is  INCORRECT.

            I came to bring the correct solution.


Let six consecutive integer numbers be  n, n+1, n+2, n+3, n+4, n+5.


The condition says


    n+1 = 11m    (1)       (The number in position two is a multiple of 11)

    n+5 = 10k    (2)        (the number in position six is a multiple of 10)


From equation (2), subtract equation (1), You will get

     4  = 10k - 11m.


Find two numbers that differ in 4 units;  the greater is multiple of 10; the smaller is a multiple of 11.


Thinking 7 seconds (or less), you will guess such numbers: they are  70 and 66.


Therefore, k= 7;  m= 6.


Thus the second number in the sequence is  11*6 = 66;  
     the sixth  number in the sequence is  10*7 = 70.


The consecutive six numbers are


    the indexes :     1     2     3     4     5     6

    the numbers :     65   66    67    68    69    70      ANSWER

Solved.

-------------

Such a sequence   // (a solution)  is  NOT  UNIQUE.

There are other similar solutions  (sequences)  with the same property.


//////////////


For your safety,  ignore the post by  @CubeyThePenguin,  since his solution is  WRONG.